OpenStudy (anonymous):

Prove: sin^2 x + tan^2 x + cot^2 x = sec^2 x

mathslover (mathslover):

@k.rajabhishek , try to get the equation only in the form of \(\large \sin^2 x\) and \(\large \cos^2 x\) : \(\large \sin^2 x + \color{blue}{\cfrac{\sin^2 x}{\cos^2 x} } + \color{red}{ \cfrac{\cos^2 x}{\sin^2 x} } = \sec^2 x \)

OpenStudy (anonymous):

@mathslover I had tried till this step.. now what to do after this?

terenzreignz (terenzreignz):

Somehow, I doubt this identity...

OpenStudy (loser66):

me too.

OpenStudy (anonymous):

are u sure?

OpenStudy (loser66):

if the last one is not cot ^2 but cos ^2 it turns true

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

Have you solved it?

OpenStudy (anonymous):

i tried but not able to..

terenzreignz (terenzreignz):

You won't, because this is not an identity... now, if you replace \(\cot^2 t\) with \(\cos^2t\) instead, then you'll have an identity.

OpenStudy (anonymous):

Allright Try to get a trigonometric Identity.